# InterFormer: Towards Effective Heterogeneous Interaction Learning for Click-Through Rate Prediction

### v1

@@ 5.3. Scalability Results
To validate the scalability of InterFormer, we carry out experiments on a large-scale internal dataset containing 70B samples in total, hundreds of non-sequence features and 10 sequences of length 200 to 1,000. In general, a 3-layer InterFormer achieves a 0.15% NE gain compared to the internal SOTA model with similar FLOPs and 24% Queries Per Second (QPS) gain. Together with feature scaling, the improvement on NE can be further enlarged with a 10% of Model FLOPs Utilization (MFU) on 512 GPUs attained. InterFormer shows great generalizability on a wide range of models showing promising ROI.
@@
Figure 2. Scalability of InterFormer on (a) feature scaling and (b) model scaling. InterFormer achieves better scaling curve than Internal, and promising layer-stacking performance on internal dataset.
On Model Scaling. Besides feature scaling, we evaluate how InterFormer performs when the model scale increases. As Figure 2(b) shows, scaling InterFormer from 1 to 4 layers achieves consistent NE gains, exhibiting good scaling properties. Specifically, compared to a single layer InterFormer, a two-layer InterFormer achieves a significant 0.13% NE gain, and 0.05% and 0.04% additional NE gains can be achieved when stacking 3 and 4 layers.
@@
For communication overhead reduction, the global arch DHEN with heavy parameters and relatively light computation tends to be FSDP ((Zhao et al., 2023)) communication bound in distributed training, while the sequence arch Transformer with much higher computation (FLOPs) to parameter ratio is normally computation bound. To alleviate such inefficiency, our parallel design of global and sequence arch allows the exposed communication from DHEN modules to overlap with sequence computation effectively, resulting in a 20% QPS improvement compared to the performing two modules sequentially.
@@ 5.5. Ablation Study
To better understand how different model designs contribute to the final performance, we ablate different exchanged information, including PMA tokens and recent interacted tokens, and different modules, including the gating module, PFFN and MHA.

### v2

@@ 5.3. Scalability Results
To validate the scalability of InterFormer, we carry out experiments on a large-scale internal dataset from Meta, containing 70B samples in total, hundreds of non-sequence features and 10 sequences of length 200 to 1,000. In general, a 3-layer InterFormer achieves a 0.15% NE gain compared to the internal SOTA model with similar FLOPs and 24% Queries Per Second (QPS) gain. Together with feature scaling, the improvement on NE can be further enlarged with a 10% of Model FLOPs Utilization (MFU) on 512 GPUs attained. InterFormer shows great generalizability on a wide range of models showing promising ROI.
@@
Figure 2. Scalability results on (a) feature scaling and (b) model scaling. InterFormer achieves better scaling curve than Internal, and promising layer-stacking performance.
On Model Scaling. In addition to feature scaling, we evaluate how InterFormer performs when the model scale increases. As Figure 2(b) shows, scaling InterFormer from 1 to 4 layers achieves consistent NE gains, exhibiting good scaling properties. Specifically, compared to a single layer InterFormer, a two-layer InterFormer achieves a significant 0.13% NE gain, with an additional 0.05% and 0.04% NE gain achieved by stacking 3 and 4 layers, respectively.
@@
For communication overhead reduction, the global arch DHEN with heavy parameters and relatively light computation tends to be FSDP (Zhao et al., 2023) communication-bound in distributed training, while the sequence arch Transformer with much higher computation (FLOPs) to parameter ratio is normally computation-bound. To alleviate such inefficiency, our parallel design of global and sequence arch allows the exposed communication from DHEN modules to overlap with sequence computation effectively, resulting in a 20% QPS improvement compared to the performing two modules sequentially.
@@ 5.5. Ablation Study
To better understand how different model designs contribute to the final performance, we ablate different exchanged information, including PMA tokens and recent interacted tokens, as well as different modules like the gating module, PFFN and MHA.